INAF queue observing, 2018 November 11-12

Observer: R. Carini, M. Faccini
Telescope Operator: D. Gonzal-Huerta
Support Astronomer: O. Kuhn
Instrument: LUCI1 (LS mode), LUCI2

Summary

Night duration (12\xB0 el twilight) 1:15-12:45 UT = 11h 30m

Observing time: 10h 45min (93%)

Weather time loss: h 45m (7%)

technical time loss 0h m (%)

Data Summary

At the sunset the sky is mostly clear, wind with gusts around 20 m/s, humidity about 20.5%

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibration Cresci a1, Saracco Luci1==2, luci2      
Mancini Hat-P-47 Mods1 + MOds2 dual grating, blue grating 4.8/5.3 mods1r.0004-0220, mods1b.0001-0143 mods2b.0004-0179 completed, some images are not recovered during the transit.
Moretti A6 Mods1 red grating   mods1r.0025 no good, acquisition wrong
Nisini telluric HIP16652 luci2 mos G210 HiRes, K,J,H   luci2.0032-0043  
Nisini field 4 luci2 mos G210 HiRes, K,J,H 1.18/3h luci2.0048-0068== K, J done, H to be completed, some images have airmass aout of spec  
Antoniucci V900Mon luci2 ls HKspec 8m (120 s each ima) luci2.0074-0077 filter completed
tell Antoniucci HIP33836 luci2 ls HKspec   luci2.0081-0082  
Calibration mods Mancini        
Calibration luci Nisini field4          

Details

Calibrations LUCI
  • flat+arc Cresci A1 luci1.0001-0014,luci2.0001-0014
  • Saracco, flatarc luci2 luci2.0015-

01:06 TO opens the dome. the deley is due to burst of wind.

01:17 collimation on the target.

1:40 TO closes the dome, gusts of wind up to 22 m/s

02:02 we open again
  • 02:20 problem with the collimation

02:29 start the acquisition of the target of Mancini Hat-P-47
  • acq =mods1r.0001-0003
  • mods1b.r0001-0003
*02:45 sci mods1r.0004-0220, mods1b.0001-0143 mods2b.0004-0179
  • seeing about 1.5"
  • seeing variable.
  • 03:38 seeing up to 2.0"
  • fitsflush red (mods1), fistflus blue (mods2). Some images are not recovered.
  • seeing more than 2.0"
  • 04:46 seeing 1.2"-1.8"
  • 04:53 seeinf0.9"-1.7"
  • 4:59* restart the script with mods1-red
  • seeing 1.5'-20''
  • 5:28* restart the script with mods2-blue
  • seeing 1,5" - 2,0"
  • 06:06* restart the sript with mods1-blue
  • seeing 1.1" - 1.8"
  • 06:21 elevation 87:19 over the limits for the guiding of the telescope. Now we are in tracking mode and it is possible to loose the target.
  • 06:28 we transit the meridian it seems we haven' lost data till now
  • 06:35 restart guiding. Apparently no data lost
  • 07:08 MODS1 BLue CCCD temperature of -11 C exceede -90C -detector warming up!
  • 07:09 MODS1 Blue CCD temperature is back normale range
  • 07:10 seeing 1 - 1,2"
  • 07:10 restart MOds1 red grating
  • 07:35 stop the integration

07:39 slew to Moretti program, target A6
  • acq mods1r-0221-0223
  • we abort and take an image to verify the the star is already in the slit mods1r.0024=
  • sci mods1r.0025
  • seeing about 1.0"
  • 08:00 seeing jump to 1.7" and decreases to 1.1"
  • seeing very variable up to 1.8", but the mean seeing is 1.2"

08:30 problem with the observing procedure. * We start again the acquisition, offset failed
  • we start again the acquisition error during the preset, we send again the script
  • 09:17 error again
  • 09:18 try again the preset and it failed
  • The problem dipend on the script AcQMOds that does not ignore "syncoffset" when only MODS1 or MODS2 was used, even if it should do it.

09:20 slew to De Carli program. we try the new procedure, but we are not sure that it is going to work.
  • we try to acquire the field in the mos using just a slit, as in the case of the telluric with LUCI, but it failes.
  • ERROR: Slit mask ID592104 is a multi-object mask!
  • You must perform a MOS mask alignemtn using an mms. ModsAlign aborting

9:35 switch to LUCI

09:50 telluric nisini field 4
  • acq luci2.0028-0031
  • sci luci2.0032-0043

10:28 slew to scientific target, field4
  • acq luci2.0044-0047
  • sci luci2.0048-0069 0069 not good
  • seeing about 1.3"
  • 11:06 seeing 1.4"
  • 11:21 seeing 1.5"
  • 11:32 seeing 1.4"
  • 11:34 seeing 1.8"
  • 11:42 seeing 1.4"
  • 12:02 seeing 1.2"

12:17 stop the integration, airmass 2.00, the images are good but not the last * slew to Antoniucci target V900 mon * 12;25 check pointing

12:27 send again the script
  • acq luci2.0070-0073
  • sci HKSpec luci2.0074-0077

12:45 telluric in HKspec
  • acq ==luci2.0078-0080
  • sci luci2.0081-0082

calibration MODS
  • bias Mancini mods1r.0235-0244,mods1b.0144-0153 mods2b.0181-0189
  • cal wave mods1 and mods2 mods1r.0245-249,mods1b.0154-155 mods2b.0190-0192
  • mos flat mods1r.0250-0259,mods1b0156-0162 mods2b.0193-199
  • slitless mods1r.0300-,mods1b.0163- mods2b.200-0212
  • send again mos flat mods2 mods2b.0213-

calibration LUCI flat arc Nisini field 4 H luci2.0083-0094 K luci2.0095-0107 J ==luci2.0108-
Topic revision: r3 - 15 Dec 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback